OpenStudy (anonymous):

What is sin [2tan^-1(3/2)]?

OpenStudy (anonymous):

Before proceeding further: \[\huge \color{green}{\textbf{Welcome To Openstudy...}}\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

So do you know formula for: \[2\tan^{-1}(x)\quad ?\]

OpenStudy (anonymous):

no I don't.

OpenStudy (anonymous):

So here it goes: \[2\tan^{-1}(x) = \sin^{-1}(\frac{2x}{1 + x^2})\]

OpenStudy (anonymous):

So here you have to just convert tan into sine inverse because we have one sine outside the brackets and that will give the answer..

OpenStudy (anonymous):

do I have to use that equation to find 2tan^-1(x)?

OpenStudy (anonymous):

is there any other way?

OpenStudy (anonymous):

See: you are given with 2tan^-1(x) here in your question x is 3/2. So firstly we have to convert this tan^-1 into sin^-1 to get our answer

OpenStudy (anonymous):

is there a way of doing it using triangles?

OpenStudy (anonymous):

See I do some part to explain it more to you: \[2 \tan^{-1}(x) = \sin^{-1}(\frac{2x}{1 + x^2})\] Here x = 3/4 so: \[\sin^{-1}(\frac{2 \times \frac{3}{2}}{1 + (\frac{3}{2})^2})\] Just solve this.

OpenStudy (anonymous):

May be, but I am not aware about that.., Sorry..

OpenStudy (anonymous):

And x = 3/2 there..

terenzreignz (terenzreignz):

Might I suggest a way that only involves a few identities that should be relatively simple?

OpenStudy (anonymous):

Yep, carry on.

OpenStudy (anonymous):

Thank you very much though @waterineyes, i'm just not familiar those formulas yet, and it just isn't the method I was looking for

terenzreignz (terenzreignz):

Well, guys, I should like you to recall that \[\huge \sin(2x)=2\sin(x)\cos(x)\]

terenzreignz (terenzreignz):

And now, let \(\large x =\tan^{-1}\left(\frac32\right)\)

OpenStudy (anonymous):

It is okay @seokhwanahn...

terenzreignz (terenzreignz):

Catch me so far, @seokhwanahn ?

OpenStudy (anonymous):

what about the 2 in front?

OpenStudy (anonymous):

nevermind, please continue

terenzreignz (terenzreignz):

Yeah, we fixed that...\[\huge \sin(2x)=2\sin(x)\cos(x)\] \[\large\sin\left[2\tan^{-1}\left(\frac32\right)\right]=2\sin\left[\tan^{-1}\left(\frac32\right)\right]\cos\left[\tan^{-1}\left(\frac32\right)\right]\]

terenzreignz (terenzreignz):

And now, we can do this, as you preferred, with triangles...

terenzreignz (terenzreignz):



terenzreignz (terenzreignz):

So, the tangent of this angle is 3/2, right?